So, 0St<4 6. Define f(t) = 34 Use the Laplace transform to solve th / + 2y + y = f(t), t€ (0,00) | M(0) = 0 7(0) = 0
Answer #1

flt)=o OSIE 4 +24 y+zy+y = f(+) y0)=0, Y(0)=0. Lly) +2414)+L(Y)=L(41+)) (3247) – $400) –Y09]+2[51LY)-410334414) = L(H)